write the slope intercept equation for the line that passes through (4, -9) and has slope -3/4. please show work
Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept. Slope (-3/4)
y = (-3/4)x + b Using order pair Pt(4,-9) to solve for b
-9 = (-3/4)*4 + b
-6 = b
y = (-3/4)x -6
graph of this line. Note: m < 0, line slants to the left